• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name D. Parno

    Log entry time 00:05:51 on December11,2008

    Entry number 252910

    keyword=Swing shift summary

    Shift workers:
    Huan Yao (TO)
    Diana Parno (3rd)

    Shift begins in production running. Big Bite high voltage had recently tripped without an alarm.

    Production runs*:

    4377: continued from previous shift. 0.03978 C
    4378: 0.03973 C

    HAPPEX run 30548 started.

    4379: 0.03995 C
    4380: 0.01859 C. Stopped early due to BB HV trip (w/ alarm).
    4381: 0.03971 C
    4382: 0.001507 C. Stopped for beam tune that did not happen.
    4383: 0.03947 C.
    4384: 0.03945 C.
    4385 begun; will finish on next shift.

    Total beam charge this shift: 0.258187 C.

    *Target temperature was oscillating throughout shift.
    ** Compton rates went up dramatically partway through the shift, although the beam position remained the same. Asymmetry was reduced, although this may be due to the worsening of the statistics.